# Matérn Kernels for Tunable Implicit Surface Reconstruction

### v1

@@ 5.2 Data-dependent Matérn Kernels
We leverage the Neural Kernel Field (NKF) framework introduced by Williams et al. (2022) to make Matérn kernels learnable. Specifically, NKF proposed to feed points $x,y\in\mathbb{R}^{d}$ through an input-conditioned neural network $\varphi:\mathbb{R}^{d}\longrightarrow\mathbb{R}^{q}$ before evaluating a kernel function, $k^{\varphi}(x,y;\mathcal{D})=k\left([x,\varphi(x;\mathcal{D})],[y,\varphi(y;\mathcal{D})]\right)$. We re-implemented the NKF framework since the authors did not provide code. We set $q=32$ for our experiments. Please refer to (Peng et al., 2020; Williams et al., 2022) for more details.
Sparse surface reconstruction and extreme generalization. We compare learned Matérn kernels against the arc-cosine kernel originally used within the NKF framework and NKSR (Huang et al., 2023) in a sparse setting on ShapeNet. Again, we sample $m=1,000$ surface points and corresponding normals for each shape and set $h=1$ for all Matérn kernels. Additionally, we evaluate Matérn kernel’s out-of-category generalization ability in an extreme setting, in which we train only on chairs and evaluate on the other 12 ShapeNet categories. Table 2 reports the results, quantified using intersection-over-union (IoU), Chamfer distance, and normal consistency (NC). Learned Matérn 1/2 and 3/2 outperform the arc-cosine kernel while being significantly faster to train. Training NKF on the entire ShapeNet dataset (consisting of approx. 30,000 shapes) takes about six days—almost twice as long as for Matérn kernels, which require about three days (measured on a single NVIDIA A100). Additionally, although Matérn 1/2 is not able to surpass NKSR, it comes very close (95.3 vs. 95.6 in IoU) while having a shorter training time. Regarding out-of-category generalization, we observe that Matérn 1/2 performs best, followed by Matérn 3/2, NKF, and NKSR. Please see Figure 5 (right) and Appendix G for qualitative results.
@@
Table 3: Robustness against noise. We compare Matérn kernels’ robustness against noise on a subset of the ShapeNet dataset (including about 1,700 shapes) using different noise levels, $\sigma\in\{0,0.0025,0.005\}$ and bandwidths, $h\in\{0.5,1,2\}$. Again, NKF* denotes a re-implementation of NKF. Bold marks best result, underline second best.
@@
Convergence speed. Finally, we analyze Matérn kernel’s convergence speed in comparison to the arc-cosine kernel as implemented in NKF, see inset figure on the right. As exemplified, Matérn 1/2 already converges after just 100 epochs, while the arc-cosine kernel requires twice as many epochs, also never being able to reach the same top accuracy. In general, we see that the smoother the kernel, the slower the convergence—an observation that perfectly matches our theoretical analysis in Section 4.1. Matérn 1/2 converges the fastest, followed by Matérn 3/2, 5/2, and the Gaussian kernel.
@@ 6 Conclusion
In this work, we have proposed to use the family of Matérn kernels for tunable implicit surface reconstruction, acting as a drop-in replacement of the recently employed first-order arc-cosine kernel within frameworks such as Neural Splines (Williams et al., 2021) and Neural Kernel Fields (Williams et al., 2022). We presented an in-depth theoretical analysis of Matérn kernels, analyzing its relation to Fourier feature mappings, Siren networks, as well as arc-cosine kernels. Our systematic evaluation confirmed our theoretical results and showed that Matérn kernels are significantly faster to compute and train, and, especially Matérn 1/2 and 3/2, lead to better surface reconstructions compared to the arc-cosine kernel—both, in a non-learnable as well as learnable regime. In the future, we plan to incorporate Matérn kernels into the NKSR framework to enhance robustness against noise.
@@ Acknowledgments
We thank Francis Williams for valuable discussions and his support in re-implementing the NKF framework and Lukas Meyer for his help during NKSR evaluation. This work was funded by the German Federal Ministry of Education and Research (BMBF), FKZ: 01IS22082 (IRRW). The authors are responsible for the content of this publication.
Appendix A Proof of Proposition 2
@@
Appendix C Proof of Theorem 3
@@
$w_{i}\sim p_{\nu}(\omega)=h^{-2\nu}C_{d,\nu}\left(\frac{2\nu}{h^{2}}+(2\pi\|\omega\|)^{2}\right)^{-(\nu+d/2)}$
@@
${=}k_{\nu}(x,y)$
where ()(*)( ∗ ) follows because $b$ is uniformly distributed, and ()(**)( ∗ ∗ ) since $w$ is distributed according to $p_{\nu}$ which is, in fact, the spectral density of $k_{\nu}$, see Eq. (10) of the main paper. ∎
Appendix D Proof of Corollary 5
@@
Appendix E Additional Details for ShapeNet Experiment
In this section, we provide additional details on the ShapeNet evaluation as presented in Section 5.1 of the main paper.
E.1 Implementation Details
@@
$f(x)=\sum_{x^{\prime}\in\mathcal{X}^{\prime}}\alpha_{i}k(x,x^{\prime})$
@@
$y_{i}=\begin{cases}+{\epsilon}&\text{if }x^{\prime}_{i}\in\mathcal{X}^{+},\\ -{\epsilon}&\text{if }x^{\prime}_{i}\in\mathcal{X}^{-}.\end{cases}$
$I\in\mathbb{R}^{2m\times 2m}$ denotes the identity matrix. We employ PyTorch’s built-in Cholesky solver to numerically stable solve for $\alpha$ and set ${\epsilon}=0.005$ for all experiments.
E.2 Further Qualitative Results
@@
Figure 6: More qualitative results on ShapeNet, comparing Matérn 1/2 and 3/2 for different values of $h$ to the arc-cosine kernel as implemented in Neural Splines (Williams et al., 2021).
Appendix F Further Details on the Surface Reconstruction Benchmark
F.1 Experimental Setting
For Neural Splines (NS; Williams et al. (2021)), we used the official implementation provided here: https://github.com/fwilliams/neural-splines. Similar to the original paper, we used 15,000 Nyström samples and did the same parameter sweep over the regularization parameter, $\lambda\in\{0,10^{-13},10^{-12},10^{-11},10^{-10}\}$. For the rest of the parameters (not mentioned in the paper), default values provided in the repository have been used, except for the grid size which we set to 512. We used exactly the same setting for Matérn kernels, except that we did a modest parameter sweep over the bandwidth parameter, $h\in\{0.5,1,2\}$. We utilized the SRB data from here: https://github.com/fwilliams/deep-geometric-prior.
Despite our best efforts, we were unable to reproduce the results for the NS kernel reported in the original paper. We run on four different GPUs, including NVIDIA RTX A2000, A40, A100, and V100 (which has been used by the authors of Neural Splines) and tested different configurations of the following parameters (see code): grid_size, eps, cg-max-iters, cg-stop-thresh, outer-layer-variance. In Table 1 of the main paper, we report the best result we could achieve on a V100, which is 0.19 for Chamfer and 3.19 for the Hausdorff distance; the authors of (Williams et al., 2021) reported 0.17 and 2.85 for Chamfer and Hausdorff distance, respectively. In order to achieve this, we had to lower outer-layer-variance form $10^{-3}$ per default to $10^{-5}$.
To measure runtime, we again used the same setting as in the original NS paper: 15,000 Nyström samples, and $\lambda=10^{-11}$. Due to the absence of further information about the experimental setting in the original paper, we used default values provided in the repository for the rest of the parameters. Runtime was measured on a single NVIDIA V100, similar to (Williams et al., 2021).
F.2 Additional Results
Per-object metrics are shown in Table 4, and more qualitative result in Figure 7.
Figure 7: Qualitative results on the Surface Reconstruction Benchmark, comparing Matérn 1/2 and 3/2 to Neural Splines (NS; Williams et al. (2021)).
@@ F.2 Additional Results
Table 4: Per-object quantitative results on the Surface Reconstruction Benchmark.
Appendix G Additional Results for Data-dependent Kernels
@@
Figure 8: Additional qualitative results for learned kernels on ShapeNet. We compare against Neural Kernel Fields (NKFs; Williams et al. (2022)) and Neural Kernel Surface Reconstruction (NKSR; Huang et al. (2023)).

### v2

@@ 4.3 Data-dependent Matérn Kernels
We leverage the Neural Kernel Field (NKF) framework introduced by Williams et al. (2022) to make Matérn kernels learnable. Specifically, NKF proposed to feed points $x,y\in\mathbb{R}^{d}$ through an input-conditioned neural network $\varphi:\mathbb{R}^{d}\longrightarrow\mathbb{R}^{q}$ before evaluating a kernel function, $k^{\varphi}(x,y;\mathcal{D})=k\left([x,\varphi(x;\mathcal{D})],[y,\varphi(y;\mathcal{D})]\right)$. We re-implemented the NKF framework since the authors did not provide code. We set $q=32$ for our experiments. Please see Williams et al. (2022) for details.
Sparse surface reconstruction and extreme generalization. We compare learned Matérn kernels against the original NKF framework (with the arc-cosine kernel) and NKSR (Huang et al., 2023) in a sparse setting on ShapeNet. Again, we sample $m=1,000$ surface points and corresponding normals for each shape and set $h=1$ for all Matérn kernels. Additionally, we evaluate Matérn kernels’ out-of-category generalization ability in an extreme setting, in which we train only on chairs and evaluate on the other 12 ShapeNet categories. Table 4 reports the results, quantified using intersection-over-union (IoU), Chamfer distance, and normal consistency (NC). Learned Matérn 1/2 and 3/2 outperform the arc-cosine kernel while being significantly faster to train. Training NKF on the entire ShapeNet dataset (consisting of approx. 30,000 shapes) takes about six days—almost twice as long as for Matérn kernels, which require about three days (measured on a single NVIDIA A100). Additionally, although Matérn 1/2 is not able to surpass NKSR, it comes very close (95.3 vs. 95.6 in IoU) while having a shorter training time. Regarding out-of-category generalization, we observe that Matérn 1/2 performs best, followed by Matérn 3/2, NKF, and NKSR. Please see Figure 6 (right) and Appendix I for qualitative results.
@@
Table 5: Robustness against noise. We compare Matérn kernels’ robustness against noise on a subset of the ShapeNet dataset using different noise levels, $\sigma\in\{0,0.0025,0.005\}$ and bandwidths, $h\in\{0.5,1,2\}$. Bold marks best result, underline second best.
@@
Convergence speed. Finally, we analyze Matérn kernels’ convergence speed in comparison to the arc-cosine kernel as implemented in NKF, see inset figure on the right. As exemplified, Matérn 1/2 already converges after just 100 epochs, while the arc-cosine kernel requires twice as many epochs, also never being able to reach the same top accuracy. In general, we see that the smoother the kernel, the slower the convergence—an observation that perfectly matches our theoretical analysis in Section 3.2. Matérn 1/2 converges the fastest, followed by Matérn 3/2, 5/2, and the Gaussian kernel.
@@ 5 Conclusion
In this work, we have proposed to use the family of Matérn kernels for implicit surface reconstruction and showed that it consistently outperforms the recently employed first-order arc-cosine kernel—both, in a non-learnable as well as learnable regime—while being significantly easier to implement, faster to compute and train, and highly scalable. We demonstrated that Matérn kernels lead to tunable surface reconstruction, and, based on a new bound of the $L_{2}$ reconstruction error, derive practical insights into how to choose their tunable bandwidth parameter. Moreover, we presented an in-depth theoretical analysis, analyzing Matérn kernels’ relation to widely used Fourier feature mappings, Siren networks, and arc-cosine kernels.
@@ Acknowledgments
We thank Francis Williams for valuable discussions and his support in re-implementing the NKF framework and Lukas Meyer for his help during NKSR evaluation. This work was funded by the German Federal Ministry of Education and Research (BMBF), FKZ: 01IS22082 (IRRW). The authors are responsible for the content of this publication. The authors gratefully acknowledge the scientific support and HPC resources provided by the Erlangen National High Performance Computing Center (NHR@FAU) of the Friedrich-Alexander-Universität Erlangen-Nürnberg (FAU) under the NHR project b112dc IRRW. NHR funding is provided by federal and Bavarian state authorities. NHR@FAU hardware is partially funded by the German Research Foundation (DFG) – 440719683.
Appendix A On the Stationarity of Matérn Kernels
As noted in Section 3.1 of the main paper, Matérn kernels are stationary: they only depend on the difference $x-y$ between two points $x,y\in\Omega$, hence being translation invariant. This is because $k(x-y)=k(x+t-(y+t))$ for $t\in\mathbb{R}^{d}$ and all stationary kernels, $k$. In contrast, the arc-cosine kernel is not translation invariant as its value depends on the absolute positions, $x,y$, see Eq. (3) of the main paper (i.e., the arc-cosine kernel can not be written as a function of $x-y$). In fact, the arc-cosine kernel is non-stationary. In the following and with the help of Figure 8, we will further discuss why stationary kernels, such as Matérn kernels, might be a better choice for kernel-based implicit surface reconstruction than non-stationary kernels, such as the arc-cosine kernel.
First, stationary kernels are naturally translation invariant. Consequently, they are independent of the absolute embedding of the input points—translating them does not change the reconstructed surface as the kernel’s value only depends on the relative distances between points. This is highly beneficial from a practical perspective since no additional steps (such as centering the point cloud) have to be taken before surface reconstruction. Being non-stationary, the arc-cosine kernel behaves differently. If we would just use the kernel as in Eq. (3) of the main paper, the reconstructed surface would vary depending on the actual location of the input points in space. This is a highly undesirable property when it comes to surface reconstruction. To make the arc-cosine kernel translation invariant, input points must be centered before reconstruction (thus eliminating the effects of translation). A centered arc-cosine kernel, denoted as $\bar{k}_{\text{AC}}$, reads:
$\bar{k}_{\text{AC}}(x,y)=\frac{\|x-\bar{c}\|\|y-\bar{c}\|}{\pi}\left(\sin\bar{\theta}+(\pi-\bar{\theta})\cos\bar{\theta}\right),\quad\text{where}\quad\bar{\theta}=\cos^{-1}\left(\frac{(x-\bar{c})^{\top}(y-\bar{c})}{\|x-\bar{c}\|\|y-\bar{c}\|}\right).$
Here, $c\in\mathbb{R}^{d}$ denotes the centroid of the input point cloud. $\bar{k}_{\text{AC}}$ is obviously translation invariant because $\|x+t-(\bar{c}+t)\|\|y+t-(\bar{c}+t)\|=\|x-\bar{c}\|\|y-\bar{c}\|$ and $(x+t-(\bar{c}+t))^{\top}(y+t-(\bar{c}+t))=(x-\bar{c})^{\top}(y-\bar{c})$ for all $t\in\mathbb{R}^{d}$. This follows because translating the input points also shifts its centroid. In conclusion, while stationary Matérn kernels can be directly used for kernel-based surface reconstruction, additional pre-processing steps are necessary for the arc-cosine kernel.
Figure 8: Matérn kernels are translation invariant and lead to locally consistent surface reconstructions. Here, we show that two identical chairs in a scene are reconstructed differently depending on the absolute position of the input points when using the non-stationary arc-cosine kernel (especially noticeable at the chair’s legs). In contrast, stationary Matérn kernels lead to locally consistent surface reconstructions, being independent of the actual embedding (i.e., both chairs have the exact same reconstructed surface).
Second, stationary kernels lead to locally consistent surface reconstructions. Since a stationary kernel’s value only depends on the relative positions, objects in a scene (or parts of objects) with similar geometric properties will be reconstructed consistently as long as the relative distances between points on each instance (or part) remain the same. This is a crucial and very desirable property for surface reconstruction which is not shared by the arc-cosine kernel (because it is non-stationary).
Appendix B Proof of Proposition 2
@@
Appendix D Proof of Theorem 3
@@
$w_{i}\sim p_{\nu}(\omega)=h^{-2\nu}C_{d,\nu}\left(\frac{2\nu}{h^{2}}+(2\pi\|\omega\|)^{2}\right)^{-(\nu+d/2)}$ (13)
@@
${=}k_{\nu}(x,y).$
Here, ()(*)( ∗ ) follows because $b$ is uniformly distributed, and ()(**)( ∗ ∗ ) since $w$ is distributed according to $p_{\nu}$ which is, in fact, the spectral density of $k_{\nu}$, see, e.g., Eq. (10) in Kanagawa et al. (2018). ∎
In fact, Theorem 3 can be generalized in that the NTK of a Siren is generally any stationary kernel:
Remark 7.
The NTK of a Siren can be associated with any stationary kernel if the bottom-layer weights in Eq. (13) are initialized according to the corresponding kernel’s spectral density.
This is a direct consequence of Bochner’s theorem.
Appendix E Proof of Corollary 5
@@
Appendix F Additional Details for ShapeNet Experiment
In this section, we provide implementation details and additional qualitative results on the ShapeNet experiment as presented in Section 4.1 of the main paper.
F.1 Implementation Details
@@
$f(x)=\sum_{i=1}^{2m}\alpha_{i}k(x,x_{i}^{\prime})$
@@
$y_{i}=\begin{cases}+{\epsilon}&\text{if }x^{\prime}_{i}\in\mathcal{X}^{+},\\ -{\epsilon}&\text{if }x^{\prime}_{i}\in\mathcal{X}^{-},\end{cases}\quad\text{for }i=1,2,\dots,2m.$
Here, $I\in\mathbb{R}^{2m\times 2m}$ denotes the identity matrix. We employ PyTorch’s built-in Cholesky solver to numerically stable solve for $\alpha$ and set ${\epsilon}=0.005$ for all experiments.
F.2 Further Qualitative Results
@@
Figure 9: Additional qualitative results on ShapeNet for non-learnable kernels. We compare the arc-cosine kernel against Matérn 1/2 and 3/2 for different values of $h$.
Appendix G Further Details on the Surface Reconstruction Benchmark
This section details the experimental setting that has been used to run Neural Splines (NS; Williams et al. (2021)) on SRB in Section 4.1 of the main paper. Moreover, we present per-object metrics and additional qualitative results.
G.1 Experimental Setting
For NS, we used the official implementation provided here: https://github.com/fwilliams/neural-splines. Similar to the original paper, we used 15,000 Nyström samples and did the same parameter sweep over the regularization parameter, $\lambda\in\{0,10^{-13},10^{-12},10^{-11},10^{-10}\}$. For the rest of the parameters (not mentioned in the paper), default values provided in the repository have been used, except for the grid size which we set to 512. We used exactly the same setting for Matérn kernels, except that we did a modest parameter sweep over the bandwidth parameter, $h\in\{0.5,1,2\}$. We utilized the SRB data from here: https://github.com/fwilliams/deep-geometric-prior.
Despite our best efforts, we were unable to reproduce the results for the NS kernel reported in the original paper. We run on four different GPUs, including NVIDIA RTX A2000, A40, A100, and V100 (which has been used by the authors of Neural Splines), and tested different configurations of the following parameters (see code): grid_size, eps, cg-max-iters, cg-stop-thresh, outer-layer-variance. In Table 1 of the main paper, we report the best result we could achieve on a V100, which is 0.19 for Chamfer and 3.19 for the Hausdorff distance; Williams et al. (2021) reported 0.17 and 2.85 for Chamfer and Hausdorff distance, respectively. In order to achieve this, we had to lower outer-layer-variance form $10^{-3}$ per default to $10^{-5}$.
To measure runtime, we again used the same setting as in the original NS paper: 15,000 Nyström samples, and $\lambda=10^{-11}$. Due to the absence of further information about the experimental setting in the original paper, we used default values provided in the repository for the rest of the parameters. Runtime was measured on a single NVIDIA V100, similar to Williams et al. (2021).
G.2 Additional Results
Per-object metrics are shown in Table 6, and more qualitative results in Figure 10.
@@ G.2 Additional Results
SAP 0.34 8.83 0.22 3.09 0.17 3.30 0.18 5.54 0.13 3.49
DiGS 0.28 5.71 0.21 5.02 0.15 2.13 0.16 3.81 0.12 1.10
VisCo 0.21 3.00 0.26 4.06 0.15 2.22 0.17 4.40 0.12 1.06
OG-INR 0.29 7.56 0.23 2.89 0.17 2.68 0.19 5.01 0.13 2.14
Table 6: Per-object quantitative results on SRB. ”NS*” denotes the best result we could achieve by running the official implementation of NS, see Section G.1. Bold marks best result, underline second best.
Figure 10: Additional qualitative results on SRB for non-learnable kernels. We compare Neural Splines (NS; Williams et al. (2021)) to Matérn 1/2 and 3/2.
Appendix H Additional Details For Texture Reconstruction Experiment
This section provides implementation details and further information about the evaluation protocol used for the texture reconstruction experiment presented in Section 4.2 of the main paper.
H.1 Implementation Details
Recall from Section 4.2 of the main paper that we are given a dataset $\mathcal{D}^{\prime}=\mathcal{D}\times\{c_{1},c_{2},\dots,c_{m}\}\subset\Omega\times\mathbb{R}^{d}\times\mathbb{R}^{3}$ with per-point RGB colors $\{c_{1},c_{2},\dots,c_{m}\}$. In order to also reconstruct texture, we are seeking a function $f^{\prime}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{4}$, predicting signed distances as well as RGB values. Using notation introduced in Section F.1, we compute $\alpha^{\prime}\in\mathbb{R}^{2m\times 4}$ as
$\alpha^{\prime}=(K+\lambda I)^{-1}y^{\prime},\quad\text{where}\quad y^{\prime}:=[y,\bar{c}]\in\mathbb{R}^{2m\times 4}$
and $\bar{c}:=[c,c]^{\top}$ with $(c_{1},c_{2},\dots,c_{m})\in\mathbb{R}^{m\times 3}$. Then, we predict the function $f^{\prime}$ as
$f^{\prime}(x)=\sum_{i=1}^{2m}\alpha_{i}^{\prime}k(x,x_{i}^{\prime}),$
where $\alpha_{i}^{\prime}\in\mathbb{R}^{4}$ denotes the $i$-th row of $\alpha^{\prime}$. Clearly, $f^{\prime}(x)=(f_{1}^{\prime}(x),f_{2}^{\prime}(x),f_{3}^{\prime}(x),f_{4}^{\prime}(x))\in\mathbb{R}^{4}$ with $f^{\prime}_{1}$ denoting the singed distance at $x$, and $f^{\prime}_{2},f^{\prime}_{3},f^{\prime}_{4}$ are the component functions that represent RGB colors at a position $x$.
To extract a textured surface mesh from the predicted four-dimensional volume, we use a two-stage process. First, we apply Marching Cubes (Lorensen & Cline, 1987) to $f^{\prime}_{1}$, and then trilinearily interpolate RGB colors at the previously predicted surface points.
H.2 Evaluation Protocol
Our quantitative evaluation on the Google Scanned Objects (GSO; Downs et al. (2022)) dataset closely follows Mitchel et al. (2024), in which authors compute image-based evaluation metrics on the texture atlas to quantify reconstruction ability. Specifically, we first sample $m$ surface points along with corresponding normals and per-point RGB colors from shapes in GSO. Next, we solve the KRR problem in Eq. (1) as always. Then, instead of predicting RGB colors on the estimated geometry (which generally differs between various kernel functions), we predict (i.e., trilinearily interpolate) colors on the surface points of the original mesh. This allows us to leverage the existing UV parametrization of the original mesh to generate a texture atlas with colors predicted from KRR. Once the texture atlas at hand, we compare it to the original texture atlas using PSNR and LPIPS.
We used ten objects from GSO with different complexity. The following assets have been used:
  • ACE_Coffee_Mug_Kristen_16_oz_cup
  • Animal_Planet_Foam_2Headed_Dragon
  • Jansport_School_Backpack_Blue_Streak
  • Now_Designs_Bowl_Akita_Black
  • Predito_LZ_TRX_FG_W
  • Reebok_ZIGTECH_SHARK_MAYHEM360
  • Razer_Kraken_Pro_headset_Full_size_Black
  • RJ_Rabbit_Easter_Basket_Blue
  • Schleich_Lion_Action_Figure
  • Weisshai_Great_White_Shark
They can be downloaded from: https://app.gazebosim.org/dashboard.
Appendix I Additional Results for Data-dependent Kernels
@@
Figure 11: Additional qualitative results on ShapeNet for learned kernels. We compare Neural Kernel Fields (NKFs; Williams et al. (2022)) and Neural Kernel Surface Reconstruction (NKSR; Huang et al. (2023)) against learnable Matérn 1/2 and 3/2.